+ cons tant retuens to scate? ren, rn St, scale . Show prove. is or inurlaing decreasing
Answer #1

F(K,L)=K^2+L^2

t>0

F(tK,tL)=(tK)^2+(tL)^2=t^2K^2+t^2L^2=t^2(K^2+L^2)=t^2F(K,L)

Thus scaling the inputs t times scales the output greater than t so the production function exhibits increasing returns to scale.